what is 50% of 50%. Write your answer as:
1.A percentage
2.a decimal
3.A common fraction
1
Expert's answer
2019-05-03T12:08:34-0400

1. 0.5*0.5=0.25=25%

2. 0.5*0.5=0.25

3. 0.5*0.5=0.25=25/100=1/4